# Expressivity of Neural Networks with Random Weights and Learned Biases

### v1

@@ Abstract
Landmark universal function approximation results for neural networks with trained weights and biases provided impetus for the ubiquitous use of neural networks as learning models in Artificial Intelligence (AI) and neuroscience. Recent work has pushed the bounds of universal approximation by showing that arbitrary functions can similarly be learned by tuning smaller subsets of parameters, for example the output weights, within randomly initialized networks. Motivated by the fact that biases can be interpreted as biologically plausible mechanisms for adjusting unit outputs in neural networks, such as tonic inputs or activation thresholds, we investigate the expressivity of neural networks with random weights where only biases are optimized. We provide theoretical and numerical evidence demonstrating that feedforward neural networks with fixed random weights can be trained to perform multiple tasks by learning biases only. We further show that an equivalent result holds for recurrent neural networks predicting dynamical system trajectories. Our results are relevant to neuroscience, where they demonstrate the potential for behaviourally relevant changes in dynamics without modifying synaptic weights, as well as for AI, where they shed light on multi-task methods such as bias fine-tuning and unit masking.
@@ 1.1 Context and Motivation
The universal approximation theorems Hornik et al. (1989); Funahashi (1989); Hornik (1991) of the late 1900s highlighted the expressivity of neural network models, i.e. their ability to approximate or express a broad class of functions through tuning of weights and biases, heralding the central role neural networks play in Machine Learning (ML) and neuroscience today. Since these foundational studies, a rich literature has explored the limits of the expressivity of neural networks by finding smaller parameter subsets whose tuning still results in the approximation of wide classes of functions or dynamics. Prior work has explored approximation capabilities of feed-forward (FFNs) and Recurrent Neural Networks (RNNs) where only the output weights are trained Rosenblatt et al. (1962); Rahimi and Recht (2008); Ding et al. (2014); Neufeld and Schmocker (2023); Jaeger and Haas (2004); Sussillo and Abbott (2009); Gonon et al. (2023); Hart et al. (2021), and deep FFNs where only subsets of parameters Rosenfeld and Tsotsos (2019), normalization parameters Burkholz (2023); Giannou et al. (2023), or binary masks–either over units or parameters–are trained Malach et al. (2020). Recently, a study has also explored the approximation abilities of transformers where only context is tuned Petrov et al. (2024). Here, motivated by recent insights from neuroscience, we initiate the investigation of expressivity in both FFNs and RNNs where only biases are trained, an avenue that remains largely unexplored.
Learned biases are of fundamental relevance to neuroscience and ML. In the latter domain, recent work has highlighted the optimization or the careful selection of bias vectors: fine-tuning of biases for multi-task learning Zaken et al. (2021) and the context tokens used for in-context learning Von Oswald et al. (2023) can all be viewed as methods of carefully setting the bias in a model where other connectivity parameters are fixed in order to perform a task. In neuroscience, single neurons employ diverse mechanisms to adjust their response to inputs, beyond synaptic plasticity. Importantly, some of these mechanisms can be construed as manipulating the firing onset of neurons which, in a firing rate model, is represented by the bias. Among these, we have shunting inhibition Holt and Koch (1997), threshold adaptation Azouz and Gray (2000), and a host of other mechanisms that participate in shaping the input-output transfer function of neurons (reviewed in Ferguson and Cardin (2020), see also Mazzucato et al. (2019)). Experimental evidence also suggests that bias-related signals play a role in learning Zhang and Linden (2003); Sehgal et al. (2013) but, despite this, most work modelling learning in neuroscience has focused on synapses.
@@
If tuning the biases of a neural network will only span a reduced set of functions, or output dynamics, then this would solidify the role of synaptic plasticity as the critical component in biological and artificial learning. Conversely, if one can express arbitrary dynamics solely by changing biases, this would motivate deeper investigation of when and how non-synaptic plasticity mechanisms might shoulder some of the effort of learning. In this paper we address the question of the expressivity of bias learning. In a regime where all weight parameters are randomly initialized and frozen, and only hidden layer biases are optimized, we provide theoretical guarantees demonstrating that:
@@
    feed-forward neural networks with wide hidden layers are universal function approximators with high probability;
@@
    RNNs with wide hidden layers can arbitrarily approximate finite-time trajectories from smooth dynamical systems with high probability.
We provide empirical support for, and a deeper interrogation of, these theoretical findings with an array of numerical experiments.
@@ 1.2 Related Works
Neuroscience. Evidence from experimental and theoretical neuroscience shows that cortical circuits flexibly regulate their dynamical phases and toggle between different regimes solely by changing the distribution of input biases to a circuit. Bias-induced modulations can explain the context-dependent effects of expectation Mazzucato et al. (2019), movements Wyrick and Mazzucato (2021) and arousal Papadopoulos et al. (2024) on sensory processing across modalities and brain areas. Changes in the bias distribution modulate the gain of the input-output neuronal transfer function Mazzucato et al. (2019). Moreover, two bias-related biological mechanisms, neural firing threshold and network inputs, have also been noted for their ability to shape network dynamics: theoretical work shows that threshold heterogeneity can improve network capacity Gast et al. (2024, 2023), while a mix of theory and experimental evidence suggests a role for inputs from upstream brain areas in reconfiguring circuit dynamics on fast timescales Perich et al. (2018); Remington et al. (2018). Within the RNN reservoir computing approach to modelling in neuroscience, where recurrent weights are random and fixed, bias modulations can toggle between multiple phases (including fixed point, chaos, and multistable regimes) and, strikingly, enable RNN multi-tasking in the absence of any parameter optimization Ogawa et al. (2023). A repertoire of dynamical motifs can also be generated in RNN reservoirs with dynamic feedback loops Logiaco et al. (2021). However, in all these studies input biases or other parameters were either i.i.d. generated or set to specific values, but never learned. The role of plastic biases has not been investigated yet and will be the main focus of our efforts below.
@@
Machine Learning. We mention primarily two bodies of literature in ML/AI that are related to this study. Many efforts have explored neural networks that are trained to quickly meta-learn new tasks via dynamics in activation space alone, without any adaptation of weights (see e.g. Feldkamp et al. (1997); Klos et al. (2020); Cotter and Conwell (1990, 1991)). Like our work, this thread of research proposes a mechanism by which a network might adapt to any new task without changing weights. However, prior work differs form the current study in that it requires an initial meta-training of all parameters in a network, weights included, before operating in the "fast learning" regime where network variables maintain context information that allow the networks to rapidly adapt to new tasks. In some cases these context variables can be thought of as biases Cotter and Conwell (1991). The closest ML contributions to our work are on the topic of masking–particularly the Strong Lottery Ticket Hypothesis (SLTH). This hypothesis conjectured that a given desired parameterization of a neural network can be found as a sub-network in a larger, appropriately initialized, network Ramanujan et al. (2020). SLTH is typically formulated with respect to weights, i.e. a subnetwork is defined by deleting weights from the full network. However, a few studies have investigated SLTH when subnetworks are constructed by deleting units Malach et al. (2020). While our study is different at face value for its focus on function approximation via bias optimization, rather than finding “lottery ticket" subnetworks, a key step in our analytic derivations relies on masking in a fashion analogous to proofs of the unit masking version of SLTH. Thus, our work also provides two results that may be of interest to the SLTH theory. First an alternative proof (complementing the first proof in Malach et al. (2020)) of the SLTH for units in single-hidden layer neural networks; second, a first proof of the SLTH over units for RNNs (see Section §2 for more details). Flavours of the lottery ticket hypothesis for RNNs has been explored empirically Yu et al. (2019); García-Arias et al. (2021); Schlake et al. (2022) but we have not encountered its proof, over weights or units, in the literature.
@@ 2 Theory Results
2.1 Feed-forward neural networks
This section studies the single-layer FFN, whose output is given by
$y_{n}(x,\theta)=\sum_{i=1}^{n}A_{:i}\phi(B_{i:}x+b),$ (1)
with $A\in\mathbb{R}^{l\times n}$, $B\in\mathbb{R}^{n\times d}$, $b\in\mathbb{R}^{n}$, and $\theta=\{B,A,b\}$. We shall investigate the approximation properties of this neural network when all the weights in $B$ and $A$ are fixed and sampled uniformly from the $n(l+d)$ dimensional centered hypercube, with half-edges of length $\gamma$, and only $b$ is tuned. We begin by outlining the activation function assumptions necessary for our theoretical results.
@@ 2.1 Feed-forward neural networks
[…]
From the universal approximation theorems of 1993 Leshno et al. (1993); Hornik (1993), a sufficient condition for $\phi$ to be a suitable activation is that it is non-polynomial. In this paper we conceptualize universal approximation as the approximation of continuous functions on compact sets with respect to an $L^{1}$ functional norm, but we remark that the literature has also studied other conditions on $h$ (for example, continuity and measurability) and other forms of convergence (for example, with respect to the $L^{\infty}$ norm). For a review of the literature, see Pinkus (1999).
@@
A suitable activation $\phi$ is referred to as a $\gamma$-parameter bounding activation if it allows for universal approximation even when each individual parameter, e.g. an element of a weight matrix or bias vector, is bounded.
@@
The proof is in Appendix §2.1 for completeness. A key subtlety of parameter-bounding is that it is a bound on individual, scalar, parameters. Thus, as a network grows in width the bias vector and weight matrix norms will still grow accordingly. This may be important, as research suggests band-limited parameters cannot universally approximate, at least for certain activation types Li et al. (2023). We leave it to future work to determine which other activations are parameter bounding.
@@
If $\phi$ is a $\gamma$-parameter bounding activation, is continuous, and if $\exists\tau\in\mathbb{R}$ such that for $x<\tau$ $\phi(x)=0$ then we say that $\phi$ is a $\gamma$-bias-learning activation.
Obviously the ReLU is a bias-learning activation. We leave the study of discontinuous functions like the Heaviside to future work. We conclude with the main result of this section. Define $p_{R}$ to be a uniform distribution on $[-R,+R]$, where $\gamma<R<\infty$.
@@
Assume that $\phi$ is $\gamma$-bias learning and, for compact $U\subset\mathbb{R}^{d}$, $h:U\to\mathbb{R}^{l}$ is continuous. Then for any degree of accuracy $\epsilon>0$ and probability of error $\delta\in(0,1)$, there exists a hidden-layer width $m\in\mathbb{N}$ and bias vector $b\in\mathbb{R}^{m}$ such that, with a probability of $1-\delta$, a neural network given by Eq.1 with each individual weight sampled from $p_{R}$ approximates $h$ with error less than $\epsilon$.
@@
Assume that $d=l$, i.e. the input and output space of $h$ has the same dimension. Then the results of Theorem 1 also hold for single-hidden-layer res-nets.
Proof Intuition: We provide intuition about the proof of Theorem 1 and its Corollary, whose details can be found in the Appendix. According to the universal approximation theorem, given a continuous function, we can find a one-hidden-layer network, $\mathcal{N}_{1}$, that is close to that function in the $L^{1}$ norm on the (compact) space of its inputs. If $\mathcal{N}_{1}$ has been constructed using $\gamma$-parameter bounding activation functions, then we know that each parameter will be on the interval $[-\gamma,\gamma]$. Next, we construct a second network, $\mathcal{N}_{2}$, to approximate $\mathcal{N}_{1}$ by randomly sampling each of its parameters, weight or bias, from $p_{R}$. For $\mathcal{N}_{2}$ to approximate $\mathcal{N}_{1}$, each parameter of $\mathcal{N}_{2}$ should fall within a tiny window of an analogous parameter in $\mathcal{N}_{1}$. This window must have half-length less than $\epsilon$ to yield the desired error bound. Without loss of generality we can assume $\epsilon<R-\gamma$. Then, if we sample parameters uniformly on $[-R,R]$ there will be a non-zero probability that a given parameter of $\mathcal{N}_{2}$ will end up within the tiny $\epsilon$-window centered at a corresponding parameter value in $\mathcal{N}_{1}$; because $\epsilon<R-\gamma$ we know that the $\epsilon$-window won’t stretch outside the distribution support. If we randomly sample a very large number of units to construct the hidden layer of $\mathcal{N}_{2}$ the probability of finding a subnetwork of $\mathcal{N}_{2}$ corresponding to $\mathcal{N}_{1}$ can be made arbitrarily close to $1$. If the activation function is bias-learning we can use biases to pick out this subnetwork by setting them appropriately smaller than the threshold given in Definition 3. We stress that this proof requires exceedingly massive hidden layer widths.

### v3

@@ Abstract
Landmark universal function approximation results for neural networks with trained weights and biases provided the impetus for the ubiquitous use of neural networks as learning models in neuroscience and Artificial Intelligence (AI). Recent work has extended these results to networks in which a smaller subset of weights (e.g., output weights) are tuned, leaving other parameters random. However, it remains an open question whether universal approximation holds when only biases are learned, despite evidence from neuroscience and AI that biases significantly shape neural responses. The current paper answers this question. We provide theoretical and numerical evidence demonstrating that feedforward neural networks with fixed random weights can approximate any continuous function on compact sets. We further show an analogous result for the approximation of dynamical systems with recurrent neural networks. Our findings are relevant to neuroscience, where they demonstrate the potential for behaviourally relevant changes in dynamics without modifying synaptic weights, as well as for AI, where they shed light on recent fine-tuning methods for large language models, like bias and prefix-based approaches.
@@ 1 Introduction
The universal approximation theorems Hornik et al. (1989); Funahashi (1989); Hornik (1991) of the late 1900s highlighted the expressivity of neural network models, i.e. their ability to approximate or express a broad class of functions through the tuning of weights and biases, heralding the central role that neural networks play in Machine Learning (ML) and neuroscience today. Since these foundational studies, a rich literature has explored the limits of this expressivity by finding smaller parameter subsets that, when optimized, can still support the approximation of wide classes of functions or dynamics. Prior work has explored the approximation capabilities of Feedforward Neural Networks (FNNs) and Recurrent Neural Networks (RNNs) where only the output weights are trained Rosenblatt et al. (1962); Rahimi & Recht (2008); Ding et al. (2014); Neufeld & Schmocker (2023); Jaeger & Haas (2004); Sussillo & Abbott (2009); Gonon et al. (2023); Hart et al. (2021), and deep FNNs where only subsets of parameters Rosenfeld & Tsotsos (2019), normalization parameters Burkholz (2023); Giannou et al. (2023), or binary masks–either over units or parameters–are trained Malach et al. (2020).
In this work, we study the expressivity of neural networks where only biases–which can be interpreted as constant inputs to neural units–are learned. While this may seem like an odd pursuit, bias-related learning is central to several active areas of research. In AI, modern sequence models such as transformers can have their outputs reshaped based on examples or instructions presented in an unchanging prefix sequence. No changes of weights are performed in this case, only distinct inputs carrying information about the context are passed to pre-trained but fixed neural network components Li & Liang (2021); Marvin et al. (2023); Garg et al. (2022); Von Oswald et al. (2023). Even more closely related, training only biases is a strategy that has been used recently Zaken et al. (2021) for more efficient fine-tuning. In neuroscience, there is growing evidence that animals can leverage inputs–via long-range projections from higher cortical areas or neuromodulatory nuclei–in order to rapidly and flexibly adapt network dynamics to multiple tasks Mazzucato et al. (2019); Ogawa et al. (2023); Perich et al. (2018); Remington et al. (2018). These tonic inputs mediating the effects of projections from other brain areas to the given, local, circuit are typically modelled as biases (see Supplementary Table 1). Consequently, input-based learning is a conceptually crucial yet poorly understood component of both modern AI systems as well as of brain functions.
@@
Quantifying the expressivity of bias learning would show the degree to which the brain or neural networks can rely on the adaptation of bias-related parameters to structure their dynamics for new tasks, thus providing critical theoretical grounding to this growing literature. If tuning the biases of a neural network will only span a reduced set of functions, or output dynamics, then this would solidify the role of synaptic plasticity as the critical component in biological and artificial learning. Conversely, if one can express arbitrary dynamics solely by changing biases, this would motivate deeper investigation of when and how non-synaptic mechanisms might shoulder some of the effort of learning. In this paper, we take a first step towards characterizing the expressivity of bias learning by studying the arguably worst-case scenario of a neural network with unstructured weights. In a regime where all weight parameters are randomly initialized and frozen, and only hidden-layer biases are optimized–which we term bias learning–we give theoretical guarantees demonstrating that:
@@
    bias-learning FNNs with wide hidden layers are universal function approximators with high probability;
@@
    bias-learning RNNs with wide hidden layers can arbitrarily approximate finite-time trajectories from smooth dynamical systems with high probability.
We further provide empirical support for, and a deeper interrogation of, these results with numerical experiments exploring multi-task learning, motor-control, and dynamical system forecasting.
@@ 1.1 Related works
Machine Learning. Many efforts have explored neural networks that are trained to quickly meta-learn new tasks via dynamics in activation space alone, without any adaptation of weights (see Feldkamp et al. (1997); Klos et al. (2020); Cotter & Conwell (1990; 1991); Hochreiter et al. (2001); Subramoney et al. (2024)). Like our work, this research proposes a mechanism by which a network might “learn” any new task without changing weights. However, prior work differs from the current study in that it requires an initial meta-training of all parameters in a network, weights included, before operating in the ”fast learning” regime where network variables maintain context information that allow the networks to rapidly adapt to new tasks. In some cases, these context variables can be thought of as biases Cotter & Conwell (1991).
From a mathematical perspective, our work is closely related to masking, particularly the Strong Lottery Ticket Hypothesis (SLTH). This hypothesis conjectures that a desired network parameterization could be found as a sub-network in a larger, appropriately-initialized, network Ramanujan et al. (2020). SLTH is typically formulated with respect to weights, i.e., a subnetwork is defined by deleting weights from the full network. However, a few studies have investigated SLTH where subnetworks are constructed by deleting units Malach et al. (2020), which we term SLTH over units. While our study is different for its focus on function approximation via bias optimization, rather than finding “lottery ticket” subnetworks, a key step in our analytic derivations relies on masking in a fashion analogous to proofs of SLTH over units. Thus, our work also provides two results that may be of interest to the SLTH theory: a novel proof of SLTH over units in single-layer FNNs, complementing the work of Malach et al. (2020) (see Connections with Malach et al. 2020 in §B.2 for details), and a first proof of SLTH for RNNs (see Section §2 for more details). Flavours of the lottery ticket hypothesis for RNNs have been explored empirically Yu et al. (2019); García-Arias et al. (2021); Schlake et al. (2022) but we have not encountered its proof, neither over weights nor units, in the literature.
Neuroscience. Changes in input biases, which mediate a change in the input-output transfer functions of neurons, can explain the context-dependent effects of expectation Mazzucato et al. (2019), movements Wyrick & Mazzucato (2021) and arousal Papadopoulos et al. (2024) on sensory processing across modalities and brain areas. Some of these effects occur by plasticity-driven changes in amygdalar projections to cortex, induced by associative learning Vincis & Fontanini (2016); Haley et al. (2020). Changes in neural firing threshold and network inputs, similar to bias modulations, were shown to shape network dynamics: threshold heterogeneity can improve network capacity Gast et al. (2024; 2023) and reconfigure circuit dynamics on fast timescales Perich et al. (2018); Remington et al. (2018). A recent study showed that, in RNNs trained to perform neuroscience tasks, learning biases via language model embedding leads to zero-shot generalization to new tasks Riveland & Pouget (2024). Within the reservoir computing approach to modelling in neuroscience, where recurrent weights are random and fixed, bias modulations can toggle between multiple phases (including fixed point, chaos, and multistable regimes) and, strikingly, enable RNN multi-tasking in the absence of any parameter optimization Ogawa et al. (2023). While slightly different than bias parameters, a repertoire of dynamical motifs can also be generated in RNN reservoirs with dynamic feedback loops Logiaco et al. (2021) and by modulating inputs in pre-trained networks Driscoll et al. (2024).
@@ 2 Theory results
2.1 Feedforward neural networks
This section studies the single-layer FNN, whose output is given by:
$y_{n}(x,\theta)=\sum_{i=1}^{n}A_{:i}\phi(B_{i:}x+b_{i}),$ (1)
with $A\in\mathbb{R}^{l\times n}$, $B\in\mathbb{R}^{n\times d}$, $b\in\mathbb{R}^{n}$, and $\theta=\{A,B,b\}$. Note that here, and throughout the paper, we adopt the notation $X_{i:}$ and $X_{:j}$ to denote the $i^{th}$ row and $j^{th}$ column, respectively, of a matrix $X$. We shall investigate the approximation properties of this neural network when all the weights in $B$ and $A$ are fixed and sampled uniformly from the $n(l+d)$-dimensional centered hypercube, where the half-edges are of length $\gamma$ and only $b$ is tuned. We begin by outlining the activation function assumptions necessary for our theoretical results.
@@ 2.1 Feedforward neural networks
The function $\phi$ is a suitable activation if, when employed in the neural network of Eq. 1, it allows for universal approximation of the following kind: for any continuous $h:U\to\mathbb{R}^{l}$ and any $\epsilon>0$, $\exists n\in\mathbb{N}$ and parameters $\theta$ s.t. $\sup_{x\in U}||h(x)-y_{n}(x,\theta)||\leq\epsilon$, where $U\subset\mathbb{R}^{d}$ is compact and |||||$\cdot$||| | ⋅ | | is the $1$-norm and will be throughout the paper.
From the universal approximation theorems of 1993 Leshno et al. (1993); Hornik (1993), a sufficient condition for $\phi$ to be a suitable activation is that it is non-polynomial. In this paper we conceptualize universal approximation as the approximation of continuous functions on compact sets with respect to an $L^{\infty}$ functional norm, but we remark that the literature has also studied other conditions on $h$ (for example measurability) and other forms of convergence. For a review of the literature, see Pinkus (1999).
@@
A suitable activation $\phi$ is referred to as a $\gamma$-parameter bounding activation if it allows for universal approximation even when each individual parameter, e.g. an element of a weight matrix or bias vector, is bounded by $\gamma$.
@@
The proof is in Appendix §B.2. A key subtlety of parameter-bounding is that it is a bound on individual, scalar, parameters. Thus, as a network grows in width the bias vector and weight matrix norms will still grow accordingly. This may be important, as research suggests band-limited parameters cannot universally approximate, at least for certain activation types Li et al. (2023). We leave it to future work to determine which other activations are parameter-bounding.
@@
If $\phi$ is a $\gamma$-parameter bounding activation, is continuous, and if $\exists\tau\in\mathbb{R}$ such that for $x<\tau$ $\phi(x)=0$, then we say that $\phi$ is a $\gamma$-bias-learning activation.
Obviously, the ReLU is a bias-learning activation. We leave the study of discontinuous functions like the Heaviside to future work. We conclude with the main result of this section. Define $p_{R}$ to be a uniform distribution on $[-R,+R]$, where $\gamma<R<\infty$.
@@
Assume that $\phi$ is $\gamma$-bias-learning and, for compact $U\subset\mathbb{R}^{d}$, $h:U\to\mathbb{R}^{l}$ is continuous. Then, for any degree of accuracy $\epsilon>0$ and probability of error $\delta\in(0,1)$, there exists a hidden-layer width $m\in\mathbb{N}$ and bias vector $b\in\mathbb{R}^{m}$ such that, with a probability of $1-\delta$, a neural network given by Eq.1 with each individual weight sampled from $p_{R}$ approximates $h$ with error less than $\epsilon$.
@@
Assume that $d=l$, i.e. the input and output spaces of the network have the same dimension. Then the results of Theorem 1 also hold for single-hidden-layer ResNets.
Proof Intuition: We provide intuition about the proof of Theorem 1 and its Corollary, whose details can be found in the Appendix (also see Fig. E.1 for visual proof intuition). According to the universal approximation theorem, given a continuous function, we can find a one-hidden-layer network, $\mathcal{N}_{1}$, that is close to that function in the $L^{\infty}$ norm on the (compact) space of its inputs. If $\mathcal{N}_{1}$ has been constructed using $\gamma$-parameter bounding activation functions, then we know that each parameter will be on the interval $[-\gamma,\gamma]$. Next, we construct a second network, $\mathcal{N}_{2}$, to approximate $\mathcal{N}_{1}$ by randomly sampling each of its parameters, weight or bias, from $p_{R}$. For $\mathcal{N}_{2}$ to approximate $\mathcal{N}_{1}$, each parameter of $\mathcal{N}_{2}$ should fall within a tiny window of an analogous parameter in $\mathcal{N}_{1}$. This window must have half-length less than $\epsilon$ to yield the desired error bound. Without loss of generality, we can assume $\epsilon<R-\gamma$. Then, if we sample parameters uniformly on $[-R,R]$, there will be a non-zero probability that a given parameter of $\mathcal{N}_{2}$ will end up within the tiny $\epsilon$-window centered at a corresponding parameter value in $\mathcal{N}_{1}$; because $\epsilon<R-\gamma$ we know that the $\epsilon$-window won’t stretch outside the distribution support. If we randomly sample a very large number of units to construct the hidden layer of $\mathcal{N}_{2}$ the probability of finding a subnetwork of $\mathcal{N}_{2}$ corresponding to $\mathcal{N}_{1}$ can be made arbitrarily close to $1$. If the activation function is bias-learning we can use biases to pick out this subnetwork by setting them appropriately smaller than the threshold given in Definition 3. We remark that this proof estimates exceedingly massive hidden-layer widths that, based on our numerical results, over-estimate the scaling of bias learning by orders of magnitude (see Remark 2 on Lemma 3). We thus view this proof not as a statement about scaling but as a statement of existence: for some sufficiently large but finite layer width, one can approximate the desired function with bias learning.